OpenStudy (anonymous):

simpilfy 7-5 (9-9y-6)

OpenStudy (anonymous):

i got y=-6 can anyone tell me if this is right

OpenStudy (goten77):

well there is no = sign so how can ya solve for y?

OpenStudy (anonymous):

i am sorry i wrote it wrong its 7-5(bracket) 9-(9y-6) (bracket)

OpenStudy (anonymous):

=

OpenStudy (goten77):

um ya mean 7-5= (9-(9y-6)) ???????????

OpenStudy (anonymous):

thats what it says simpilfy 7-5(bracket) 9-(9y-6) (bracket) =

OpenStudy (anonymous):

\[7-5 (9-(9 y-6))=45 y-68 \]

OpenStudy (anonymous):

the minus sign between the 9-(9y-6) is throwing me off

OpenStudy (anonymous):

45y - 68 was right thank you for all your help

OpenStudy (anonymous):

You're welcome.